Please answer both.
[10] #6. Circle all the FALSE statements below. a) The vector equation F-4 v x B gives the force F on a particle with charge q that is moving with velocity v in a magnetic field B. b The right-hand rules given apply to only negative charges. c) Like magnetic poles repel and unlike magnetic poles attract d) The Hall effect can be used to determine whether the majority charge carriers in a material have a positive, or, negative charge. e) In P-type semiconductor materials, the majority charge carriers looks like electrons, except that they are positive (8 #7 A heating element is made by maintaining a potential difference of i 2.0 V between the ends of a certain wire with a .52 x 10 m2 cross section area and a resistivity of 4.86 x 106 ohm.m. If the element dissipates 10.0 W, what is its length? Answer: (a) 0.01 m. (b) 0.08 m. (c) 12.4 m. () 18.4 m. (e) 22.3 m.(
Answer #1

#6 B , E

#7

power dissipate P = V^2/R  

R = rho*L/ A

P = V^2 / ( rho*L/ A  ) = 12^2 / ( 4.56*10^-6 * L /(7.52*10^-6)) = 10

form here i got L = 22.2815 m E is ur answer